# AdS3 Integrability, Tensionless Limits, and Deformations: A Review

### v3

@@ 3.1 Symmetries
The charges of various states are contained in Table 1 for $k\neq 0$ and in Table 2 for $k=0$. It is also worth mentioning that in general (by semi-classical arguments as well as by an analysis of the physical poles of the S matrix, see below), we expect new representations to emerge as bound states of the above representations. Specifically, left and left representations would make bound states, as would right with right. In total, for reasons which we will describe in the subsection below, we expect to have two massless representations and $k-1$ massive representation ($k-2$ of which are bound states) Frolov:2023lwd ; Frolov:2024pkz . [^17] In particular, the right representation with $\mu=-1$ can be related to that with $\mu=k-1$, which is a bound state of $(k-1)$ left excitations. If $k=0$, we expect infinitely many bound states Borsato:2013hoa .
@@
The parameter $\alpha$, with $10<\alpha<$ 1, encodes the relative radii of the two spheres. More specifically, if $R_{AdS}$ is the $AdS_{3}$ radius and $R_{1}$, $R_{2}$ are the radii of the two spheres we have
@@
so that $\alpha=\tfrac{1}{2}$ is the special case where the two spheres are identical, and $\mathfrak{d}(2,1;\tfrac{1}{2})\cong\mathfrak{osp}(4|2)$. The other notable limit is $\alpha\to 0$ (or equivalently $\alpha\to 1$) where one of the two spheres becomes flat, $S^{3}\to\mathbb{R}^{3}$, and the superalgebra contracts. This family of backgrounds, for any $10<\alpha<$ 1, is integrable Babichenko:2009dk . This remains true even if we introduce a Kalb-Ramond field in the NLSM action Cagnazzo:2012se . One major difference with the case at hand is that the most supersymmetric BMN solution is 1/4-BPS for this background, rather than 1/2-BPS; this is a geodesic that run time $t$ in $AdS_{3}$ and along great circles parametrised by $\varphi_{1}$ and $\varphi_{2}$ along either sphere (with a specific velocity for each space). As a result the algebra $\mathcal{A}$ of eq. (36) contains at most one copy of $\mathfrak{su}(1|1)^{\oplus 2}/\mathfrak{u}(1)$, centrally extended by $\mathbf{C},\mathbf{C}^{\dagger}$. As a result, short representations of $\mathcal{A}$ are two-dimensional (one boson and one fermion) rather than four-dimensional, and there are eight, rather than four, distinct representations. The near-BMN bosonic Hamiltonian still takes the form (24) but now the allowed values of $\mu$ are
@@ 3.2 S matrix
Hence, from the point of view of the representation structure, if one restricts the momentum to a fundamental region of size $2\pi$ one must consider all values of $\mu$. Viceversa, if we allow for the momentum to take any real value, one may restricts to $\mu=0,1,\dots,k$ [^20] The reason why we are counting $k+1$, rather than $k$, particles is that we expect two “massless” representations, one of which can be taken to be $\mu=0$, with the other being $\mu=k$. Strictly speaking, this is a property of the representation (and hence of the matrix part of the S matrix), and it is non-trivial to demonstrate that it is possible to construct dressing factors compatible with the requirement (47). Notice that instead for pure-RR backgrounds ($k=0$) the dispersion relation is $2\pi$-periodic, and has no (quasi)-periodicity in $\mu$. Indeed, the bootstrap procedure allows to construct representations with any $\mu\in\mathbb{Z}$, while $p$ can be restricted to a fundamental region.
@@ 4.1 The case of $k=0$ (pure-RR background)
Let us take a look at the field content of this theory. The vector multiplet contains scalar fields of the type $\Phi^{\alpha\dot{\alpha}}$, in the bifundamental representation of the $\mathfrak{su}(2)_{\mbox{\tiny L}}\oplus\mathfrak{su}(2)_{\mbox{\tiny R}}$ R-symmetry, with dimension $(1/2,1/2)$ under $(L^{0}_{\mbox{\tiny L}},L^{0}_{\mbox{\tiny R}})$ (as well as their superpartners, which we will not discuss here, see OhlssonSax:2014jtq ). Notice that this dimension is unusual for scalars in a two-dimensional (S)CFT; it is a result of having integrated out the fundamental hypermultiplet fields, which yields a non-standard (non-local) kinetic term to the $\Phi$s. One of these fields, say $\Phi^{+\dot{+}}$ can be identified with the half-BPS vacuum of the light-cone gauge; the sphere bosons $y$, $\bar{y}$ correspond to R-symmetry currents relating $\Phi^{+\dot{+}}$ to $\Phi^{-\dot{+}}$ and $\Phi^{+\dot{+}}$ to $\Phi^{+\dot{-}}$, respectively. This is quite similar to what happens in $\mathcal{N}=4$ SYM. Conversely, the adjoint hyper contains scalars of the type $T^{a\dot{a}}$, charged under $\mathfrak{su}(2)_{\bullet}\oplus\mathfrak{su}(2)_{\circ}$, which have the usual scalar field end hence have dimension $(0,0)$ under $(L^{0}_{\mbox{\tiny L}},L^{0}_{\mbox{\tiny R}})$ (like a “regular” scalar in two dimensions). They correspond to excitations from the $T^{4}$ part of the string geometry. This model seems to have nice features, including integrability OhlssonSax:2014jtq . However, starting from a reference BPS state such as
@@
[…]
@@ 4.2 The case of $k=1$ (symmetric-product orbifold CFT)
as it can be seen e.g. by expanding the stress energy tensor in terms of the fractionary modes of the fields (62); the difference between even and odd sectors is due to Fermions. In fact, in the NS sectors the chiral Fermions have modes indexed by $(-\tfrac{1}{2}-\tfrac{n}{w})$, cf. (62); similarly, the anti-chiral modes are indexed by $(-\tfrac{1}{2}-\tfrac{\tilde{n}}{w})$. Hence, the modes with $n>-w/2$ and $\tilde{n}>-w/2$ are always creation modes, while those with $n<-w/2$ and $\tilde{n}<-w/2$ annihilate $|\sigma_{w}\rangle$; when $w$ is even, the modes with $n=w/2$ and $\tilde{n}=w/2$ give rise to a Clifford module.
@@
This is of course expected from the form of the centrally extended algebra which we discussed above — which mandates $\{\mathbf{Q}_{\mbox{\tiny L}},\mathbf{Q}_{\mbox{\tiny R}}\}\neq 0$ for $h\neq 0$ — and in particular from Table (1). In fact, comparing the above algebraic description at $k=1$ and small $h$ [^28] It is worth keeping in mind that the analysis is drastically different at $k=0$ and small $h$, as discussed in the previous subsetion. with this conformal perturbation theory set-up, we find that (as expected) the RR parameter $h$ and the conformal perturbation theory parameter $g$ should be related as
@@
Notice that the twist sector of the two operators change from $w$ to $(w+1)$; this is both required by the orbifold construction and by the length-changing nature of the algebra, see eq. (35). This computation was first discussed in Gava:2002xb (well before the relation between these centrally-extended algebras and integrability came to the fore Beisert:2005tm ) and recently revisited in Gaberdiel:2023lco where it was argued that it gives a non-zero result (albeit not quite correctly). [^30] As detailed in Fabri:2025rok there is an inconsistency in that computation: in particular, the level-matching / physical state condition is not relaxed in an appropriate way, and as a result the central extensions $\mathbf{C}$, $\mathbf{C}^{\dagger}$ actually vanish in the algebra (34) when it is computed from the orbifold. Due to a separate issue in the derivation, the authors of Gaberdiel:2023lco miss this fact and claim instead to find a non-vanishing result. In particular, such a computation would allow to fix the proportionality constant in (79) to $\xi=1$. Under this assumption, the precise identification of the excitations from the orbifold with the integrability ones was performed in Frolov:2023pjw , where it was also shown that the algebra reproduces the one discussed of Lloyd:2014bsa and therefore yields their S matrix for the modes with $\mu=0$. This also fits with the general argument discussed above that, for $k=1$, we only expect excitations with $\mu=0$: recall that the dispersion relation is invariant under shifts of $\mu\to\mu-kn$ and $p\to p+2\pi n$, with $n\in\mathbb{Z}$, cf. (47). In fact, this is not just an algebraic property of the representations, but it holds for the whole S matrix (including the dressing factor Frolov:2024pkz ; Frolov:2025uwz ). Because of this, if we allow the real part of $p$ to take any value, we can identify the species of particles by $\mu~{}\text{mod}\,k$, rather than by $\mu$.
@@ 5.1 Overview of integrable deformations
TsT transformations. Perhaps the simplest integrable deformations are T-duality-shift-T-duality (TsT) transformations, see Lunin:2005jy ; Frolov:2005dj ; Alday:2005ww . These transformations require the presence of two abelian isometries in the background. For concreteness we will assume these to be realised as shifts in the two coordinates $x_{1}$ and $x_{2}$. The deformed background is obtained through a T-duality $x_{1}\rightarrow\tilde{x}_{1}$, where $\tilde{x}_{1}$ denotes the T-dual coordinate, followed by a shift $x_{2}\rightarrow x_{2}+s\tilde{x}_{1}$, and finally a T-duality back $\tilde{x}_{1}\rightarrow\tilde{\tilde{x}}_{1}$. Here $s\in\mathbb{R}$ denotes the continuous and real deformation parameter. What makes TsT deformations particularly simple is that they can be re-absorbed in the boundary conditions of the fields Frolov:2005dj ; Alday:2005ww . In the particular case where the TsT is taken (partially or completely) along one of the directions used for the lightcone gauge fixing, the interpretation of the twist of the boundary conditions is more subtle, and can be related to a “$T\bar{T}$” deformation of the gauge fixed model, see Sfondrini:2019smd ; Apolo:2019zai ; Idiab:2024bwr .
@@
for all elements $X,Y\in\mathfrak{g}$and with $c\in\mathbb{C}$ These two conditions ensure the classical integrability of the model Klimcik:2008eq . The first condition is the antisymmetry of the operator with respect to the ad-invariant bilinear form Tr, while the second is the classical Yang-Baxter equation. This statement can be made more transparent by defining $r\in\mathfrak{g}\wedge\mathfrak{g}$ and the Casimir $t$ through
@@ 5.2 Homogeneous and inhomogeneous Yang-Baxter deformations
$[H_{j},E^{\pm}_{k}]=\pm A_{jk}E^{\pm}_{k}\,,\qquad[E^{+}_{j},E^{-}_{k}]=\delta_{jk}\frac{q^{H_{k}}-q^{-H_{k}}}{q-q^{-1}}\,.$ (107)
@@ 6.1 Deformed background
    $\Phi$ $=0\,,\qquad H_{3}=\frac{\sqrt{1-\hat{q}^{2}}}{\sqrt{1+\kappa^{2}}}d\hat{B}\,,\qquad$ (129)
@@ 6.2 Deformed symmetries and S-matrix
[…]
$\Delta[\mathbf{H}_{\star}]=\mathbf{H}_{\star}\otimes 1+1\otimes\mathbf{H}_{\star}\,,\qquad$ $\Delta[\mathbf{U}]=\mathbf{U}\otimes\mathbf{U}\,,$ (138)
$\Delta[\mathbf{Q}_{\star}]=\mathbf{Q}_{\star}\otimes 1+q_{\star}^{\mathbf{H}_{\star}}\mathbf{U}^{+1}\otimes\mathbf{Q}_{\star}\,,\qquad$ $\Delta[\mathbf{S}_{\star}]=\mathbf{S}_{\star}\otimes q_{\star}^{-\mathbf{H}_{\star}}+\mathbf{U}^{-1}\otimes\mathbf{S}_{\star}\,.$
@@
$[\mathbf{H}_{\mbox{\tiny L}}]_{q_{\mbox{\tiny L}}}[\mathbf{H}_{\mbox{\tiny R}}]_{q_{\mbox{\tiny R}}}=\mathbf{P}\,\mathbf{K}\,.$ (141)
@@ 7 Elliptic deformation
$||\mathbf{x}^{(1)}||^{2}$ $=\frac{(\alpha_{2}+\alpha_{3}-\alpha_{1})}{\alpha_{2}\alpha_{3}}-||\mathbf{x}^{(4)}||^{2}\,,$ $\qquad\mathbf{x}^{(1)}\cdot\mathbf{x}^{(4)}=-\mathbf{x}^{(2)}\cdot\mathbf{x}^{(3)}\,,$ (159)
$||\mathbf{x}^{(2)}||^{2}$ $=\frac{(\alpha_{2}-\alpha_{1}-\alpha_{3})}{\alpha_{1}\alpha_{3}}+||\mathbf{x}^{(4)}||^{2}\,,$ $\qquad\mathbf{x}^{(2)}\cdot\mathbf{x}^{(4)}=+\mathbf{x}^{(1)}\cdot\mathbf{x}^{(3)}\,,$
$||\mathbf{x}^{(3)}||^{2}$ $=\frac{(\alpha_{2}+\alpha_{1}-\alpha_{3})}{\alpha_{1}\alpha_{2}}-||\mathbf{x}^{(4)}||^{2}\,,$ $\qquad\mathbf{x}^{(3)}\cdot\mathbf{x}^{(4)}=-\mathbf{x}^{(1)}\cdot\mathbf{x}^{(2)}\,.$
The classical and quantum integrability of the proposed elliptic deformation of the $AdS_{3}\times S^{3}\times T^{4}$ superstring is an open question. Insight can be obtained through computing the perturbative S-matrix including the contributions coming from the scattering with fermions, which is work in progress HRS . The fate of the broken right-acting symmetries under the elliptic deformation is not yet known, and the exact S-matrix remains to be found.

### v4

@@ 3.1 Symmetries
The charges of various states are contained in Table 1 for $k\neq 0$ and in Table 2 for $k=0$. It is also worth mentioning that in general (by semi-classical arguments as well as by an analysis of the physical poles of the S matrix, see below), we expect new representations to emerge as bound states of the above representations. Specifically, left and left representations would make bound states, as would right with right. In total, for reasons which we will describe in the subsection below, we expect to have two massless representations and k1k-1 massive representation (k2k-2 of which are bound states) Frolov:2023lwd ; Frolov:2024pkz . [^17] In particular, the right representation with $\mu=-1$ can be related to that with $\mu=k-1$, which is a bound state of (k1)(k-1) left excitations. If $k=0$, we expect infinitely many bound states Borsato:2013hoa .
@@
The parameter $\alpha$, with $0<\alpha<1$, encodes the relative radii of the two spheres. More specifically, if $R_{AdS}$ is the $AdS_{3}$ radius and $R_{1}$, $R_{2}$ are the radii of the two spheres we have
@@
so that $\alpha=\tfrac{1}{2}$ is the special case where the two spheres are identical, and $\mathfrak{d}(2,1;\tfrac{1}{2})\cong\mathfrak{osp}(4|2)$. The other notable limit is $\alpha\to 0$ (or equivalently $\alpha\to 1$) where one of the two spheres becomes flat, $S^{3}\to\mathbb{R}^{3}$, and the superalgebra contracts. This family of backgrounds, for any $0<\alpha<1$, is integrable Babichenko:2009dk . This remains true even if we introduce a Kalb-Ramond field in the NLSM action Cagnazzo:2012se . One major difference with the case at hand is that the most supersymmetric BMN solution is 1/4-BPS for this background, rather than 1/2-BPS; this is a geodesic that run time $t$ in $AdS_{3}$ and along great circles parametrised by $\varphi_{1}$ and $\varphi_{2}$ along either sphere (with a specific velocity for each space). As a result the algebra $\mathcal{A}$ of eq. (36) contains at most one copy of $\mathfrak{su}(1|1)^{\oplus 2}/\mathfrak{u}(1)$, centrally extended by $\mathbf{C},\mathbf{C}^{\dagger}$. As a result, short representations of $\mathcal{A}$ are two-dimensional (one boson and one fermion) rather than four-dimensional, and there are eight, rather than four, distinct representations. The near-BMN bosonic Hamiltonian still takes the form (24) but now the allowed values of $\mu$ are
@@ 3.2 S matrix
Hence, from the point of view of the representation structure, if one restricts the momentum to a fundamental region of size $2\pi$ one must consider all values of $\mu$. Viceversa, if we allow for the momentum to take any real value, one may restricts to $\mu=0,1,\dots,k$. [^20] The reason why we are counting $k+1$, rather than $k$, particles is that we expect two “massless” representations, one of which can be taken to be $\mu=0$, with the other being $\mu=k$. Strictly speaking, this is a property of the representation (and hence of the matrix part of the S matrix), and it is non-trivial to demonstrate that it is possible to construct dressing factors compatible with the requirement (47). Notice that instead for pure-RR backgrounds ($k=0$) the dispersion relation is $2\pi$-periodic, and has no (quasi)-periodicity in $\mu$. Indeed, the bootstrap procedure allows to construct representations with any $\mu\in\mathbb{Z}$, while $p$ can be restricted to a fundamental region.
@@ 4.1 The case of $k=0$ (pure-RR background)
[…]
@@
and considering excitations on top of it, there is the possibility of a huge amount of mixing due to the presence of the dimensionless $T^{a\dot{a}}$ fields and of their superpartners. This fact is reflected in the presence of strong wrapping corrections (not suppressed in $h\ll 1$) in the TBA. In fact, it can be checked that even at small-$h$ the wrapping effects due to TBA scale as $1/L$, i.e. that they are not suppressed at all, [^24] The modifying the S matrix appearing in the Asymptotic Bethe Ansatz itself would give corrections of order $1/L$. see Figure 3.
@@ 4.2 The case of $k=1$ (symmetric-product orbifold CFT)
as it can be seen e.g. by expanding the stress energy tensor in terms of the fractionary modes of the fields (62); the difference between even and odd sectors is due to Fermions. In fact, in the NS sectors the chiral Fermions have modes indexed by $(-\tfrac{1}{2}-\tfrac{n}{w})$, cf. (62); similarly, the anti-chiral modes are indexed by $(-\tfrac{1}{2}-\tfrac{\tilde{n}}{w})$. Hence, the modes with n>w/2n>-w/2 and $\tilde{n}>-w/2$ are always creation modes, while those with n<w/2n<-w/2 and $\tilde{n}<-w/2$ annihilate $|\sigma_{w}\rangle$; when $w$ is even, the modes with $n=w/2$ and $\tilde{n}=w/2$ give rise to a Clifford module.
@@
This is of course expected from the form of the centrally extended algebra which we discussed above — which mandates $\{\mathbf{Q}_{\mbox{\tiny L}},\mathbf{Q}_{\mbox{\tiny R}}\}\neq 0$ for $h\neq 0$ — and in particular from Table (1). In fact, comparing the above algebraic description at $k=1$ and small hh [^28] It is worth keeping in mind that the analysis is drastically different at $k=0$ and small $h$, as discussed in the previous subsetion. with this conformal perturbation theory set-up, we find that (as expected) the RR parameter $h$ and the conformal perturbation theory parameter $g$ should be related as
@@
Notice that the twist sector of the two operators change from $w$ to $(w+1)$; this is both required by the orbifold construction and by the length-changing nature of the algebra, see eq. (35). This computation was first discussed in Gava:2002xb (well before the relation between these centrally-extended algebras and integrability came to the fore Beisert:2005tm ) and recently revisited in Gaberdiel:2023lco where it was argued that it gives a non-zero result. [^30] Note that it was argued in Fabri:2025rok that there is an inconsistency in the derivation of Gaberdiel:2023lco , specifically in how the large-$w$ limit of the orbifold results is taken. According to Fabri:2025rok , the central extensions $\mathbf{C}$, $\mathbf{C}^{\dagger}$ would actually vanish in the algebra (34) when computed from the orbifold, suggesting that a different analysis of that CFT would be needed to construct the central extensions. In particular, such a computation would allow to fix the proportionality constant in (79) to $\xi=1$. Under this assumption, the precise identification of the excitations from the orbifold with the integrability ones was performed in Frolov:2023pjw , where it was also shown that the algebra reproduces the one discussed of Lloyd:2014bsa and therefore yields their S matrix for the modes with $\mu=0$. This also fits with the general argument discussed above that, for $k=1$, we only expect excitations with $\mu=0$: recall that the dispersion relation is invariant under shifts of $\mu\to\mu-kn$ and $p\to p+2\pi n$, with $n\in\mathbb{Z}$, cf. (47). In fact, this is not just an algebraic property of the representations, but it holds for the whole S matrix (including the dressing factor Frolov:2024pkz ; Frolov:2025uwz ). Because of this, if we allow the real part of $p$ to take any value, we can identify the species of particles by $\mu~\text{mod}\,k$, rather than by $\mu$.
@@ 5.1 Overview of integrable deformations
TsT transformations. Perhaps the simplest integrable deformations are T-duality-shift-T-duality (TsT) transformations, see Lunin:2005jy ; Frolov:2005dj ; Alday:2005ww . These transformations require the presence of two abelian isometries in the background. For concreteness we will assume these to be realised as shifts in the two coordinates $x_{1}$ and $x_{2}$. The deformed background is obtained through a T-duality $x_{1}\rightarrow\tilde{x}_{1}$, where $\tilde{x}_{1}$ denotes the T-dual coordinate, followed by a shift $x_{2}\rightarrow x_{2}+s\tilde{x}_{1}$, and finally a T-duality back $\tilde{x}_{1}\rightarrow\tilde{\tilde{x}}_{1}$. Here $s\in\mathbb{R}$ denotes the continuous and real deformation parameter. What makes TsT deformations particularly simple is that they can be re-absorbed in the boundary conditions of the fields Frolov:2005dj ; Alday:2005ww . In the particular case where the TsT is taken (partially or completely) along one of the directions used for the lightcone gauge fixing, the interpretation of the twist of the boundary conditions is more subtle, and can be related to a $T\bar{T}$” deformation of the gauge fixed model, see Sfondrini:2019smd ; Apolo:2019zai ; Idiab:2024bwr .
@@
for all elements $X,Y\in\mathfrak{g}$and with $c\in\mathbb{C}$ . These two conditions ensure the classical integrability of the model Klimcik:2008eq . The first condition is the antisymmetry of the operator with respect to the ad-invariant bilinear form Tr, while the second is the classical Yang-Baxter equation. This statement can be made more transparent by defining $r\in\mathfrak{g}\wedge\mathfrak{g}$ and the Casimir $t$ through
@@ 5.2 Homogeneous and inhomogeneous Yang-Baxter deformations
$H_{j},E^{\pm}_{k}]=\pm A_{jk}E^{\pm}_{k}\,,\qquad[E^{+}_{j},E^{-}_{k}]=\delta_{jk}\frac{q^{H_{k}}-q^{-H_{k}}}{q-q^{-1}}\,.$ (107)
@@ 6.1 Deformed background
    $\Phi$ $=0\,,\qquad H_{3}=\frac{\sqrt{1-\hat{q}^{2}}}{\sqrt{1+\kappa^{2}}}d\hat{B}\,$$\qquad$ (129)
@@ 6.2 Deformed symmetries and S-matrix
Let us mention that the superalgebra $\mathfrak{psu}(1|1)$ only admits one Dynkin diagram $\otimes$ (with one single fermionic simple root), and hence there is essentially only one choice of quantum group deformation. The deformed coproduct, including the braiding $\mathbf{U}$, reads (for $\star=\{\mbox{\tiny L},\mbox{\tiny R}\}$ and dropping the $a$ indices),
$\Delta[\mathbf{H}_{\star}]=\mathbf{H}_{\star}\otimes 1+1\otimes\mathbf{H}_{\star}\,$$\qquad$ $\Delta[\mathbf{U}]=\mathbf{U}\otimes\mathbf{U}\,$, (138)
$\Delta[\mathbf{Q}_{\star}]=\mathbf{Q}_{\star}\otimes 1+q_{\star}^{\mathbf{H}_{\star}}\mathbf{U}^{+1}\otimes\mathbf{Q}_{\star}\,$$\qquad$ $\Delta[\mathbf{S}_{\star}]=\mathbf{S}_{\star}\otimes q_{\star}^{-\mathbf{H}_{\star}}+\mathbf{U}^{-1}\otimes\mathbf{S}_{\star}\,.$
@@
$\mathbf{H}_{\mbox{\tiny L}}]_{q_{\mbox{\tiny L}}}[\mathbf{H}_{\mbox{\tiny R}}]_{q_{\mbox{\tiny R}}}=\mathbf{P}\,\mathbf{K}\,.$ (141)
@@ 7 Elliptic deformation
[…]
The classical and quantum integrability of the proposed elliptic deformation of the $AdS_{3}\times S^{3}\times T^{4}$ superstring is an open question. Insight can be obtained through computing the perturbative S-matrix including the contributions coming from the scattering with fermions Hoare:2025rtl . The fate of the broken right-acting symmetries under the elliptic deformation is not yet known, and the exact S-matrix remains to be found.